# Phase domain walls in coherently driven Bose-Einstein condensates

### v3

@@ Abstract
We consider coherent states of weakly interacting bosons under the conditions of external resonant excitation, with a focus on a two-dimensional polariton fluid driven by a plane electromagnetic wave near the ground state. The coherent driving breaks the U(1) symmetry explicitly, which prevents the occurrence of quantum vortices in a uniform scalar condensate. Surprisingly, a spinor (two-component) system of the same kind admits topological excitations, such as domain walls of relative phase or confined half-vortex molecules, typical of a freely evolving spinor Bose system. Opposite-phase domains arise from the spontaneous breakdown of the spin symmetry $(\mathbb{Z}_{2})$. Domain walls form with time even when the initial state of the system is uniform or completely disordered; they fall into different topological types distinguished by the total phase variation in the transverse direction.
@@ I Introduction
A coupled pair of spatially overlapping Bose-Einstein condensates (BECs) has complex topological excitations that cannot be reduced to the case of one scalar BEC. Typical examples of such systems include atomic BECs with coexistent hyperfine species [1] and two-band superconductors [2]. In one dimension, they admit the sine-Gordon kinks characterized by the $2\pi$ jump of the relative phase of order parameters [3, 4]; in two dimensions, the kink turns into a domain wall which gives birth to more complicated objects such as bound states of half-quantized vortices (HQVs) [5, 6]. The mechanism of the HQV pairing on a domain wall was found to be formally similar to quark confinement [7, 8]. The underlying mean-field theory attracts much interest owing to its broad scope and very rich physics it contains in spite of conceptual simplicity [9].
In this work, we investigate a dissipative BEC coupled to an external classical field whose frequency is close to the BEC ground-state level. Such a system is found to admit stable domain walls and HQV molecules which arise spontaneously even when the initial conditions are uniform or, conversely, disordered. The uniform BEC undergoes a spontaneous $\mathbb{Z}_{2}$ symmetry breaking that proceeds nonglobally, by way of the Kibble-Zurek mechanism [10, 11, 12], as soon as the driving field reaches a critical amplitude; in turn, a disordered system exhibits self-organization. Once dynamic equilibrium has established, a number of HQV molecules and domain walls act as elements of complex, inhomogeneous yet internally ordered states whose spatial extent is far beyond any conventional characteristic scale of a BEC.
@@ II Symmetry considerations
The last term in (3) represents the source of excitation. In the case of a one-component system (i. e., when one of $\psi_{\pm}$ is zero or $\psi_{+}$ and $\psi_{-}$ are identical), this driving term invalidates the conventional steady-state vortex solutions such that $\nabla n\neq 0$ and $\nabla\phi\neq 0$ but $(\nabla n)(\nabla\phi)=n\nabla^{2}\phi=0$; moreover, all actual steady states prove to be uniform [30]. On the other hand, the two-component system under certain conditions has a weak symmetry of the form $\psi_{+}^{\vphantom{*}}+\psi_{-}^{\vphantom{*}}=\psi_{+}^{*}+\psi_{-}^{*}$, so that quantity
$K=-\left(\sqrt{n_{+}}\sin\phi_{+}+\sqrt{n_{-}}\sin\phi_{-}\right)$ (5)
turns to zero even when both components vary in space. In this case the driving term vanishes from (3) irrespective of $f$.
The considered weak symmetry $K=0$ is equivalent to invariance of the density of energy $(\mathcal{H})$ under the transformation $\psi_{\pm}\mapsto\psi_{\pm}e^{i\chi}$ for small $\chi$. Directly from (2) one can see that $\mathcal{H}\bigl{(}\psi_{\pm}e^{i\chi}\bigr{)}-\mathcal{H}\bigl{(}\psi_{\pm}\bigr{)}$ tends to $2\chi fK$ when $\chi\to 0$. In fact, it is just a special case of Noether’s theorem that invariance under infinitesimal global phase shift yields the continuity equation with no external sources (e. g., [31]). Thus, condition $K=0$ means that not only moduli $\sqrt{n_{\pm}}$ but also phases $\phi_{\pm}$ of order parameters are partially unlocked. As a result, certain features of freely evolving BECs, such as spontaneously formed HQV molecules and moving domain walls, are no longer forbidden.
An important question is when such a symmetry is obeyed. Since the theory is spin-symmetric, Eqs. (1) always admit trivial uniform solutions with $\psi_{+}=\psi_{-}$ which have fixed phases. They, however, lose stability in some range of $f$, giving way to a doublet of the true vacuum states with $\psi_{+}^{*}=\psi_{-}^{\vphantom{*}}$ at $\gamma\to 0$. The stationary $q=0$ domain walls as well as HQV molecules represent the nonuniform states with $K=0$. The symmetry of the $q=\pm 1$ domain walls is reduced further, namely, function $K(\mathbf{r})$ becomes nonzero and antisymmetric across the wall, which manifests broken parity. Nevertheless, all domain walls are capable of a solitonic motion, because the mean value of $K$ is zero.
@@
The above considerations help us to compare our results to the previously known mechanisms of vortex formation in resonantly excited polariton systems. In all cases, the key point is to get rid of the phase locking of the driven fluid with respect to the driving optical field. Previously, this was done by making the system explicitly inhomogeneous in space or time. For instance, vortices can be “injected” immediately by means of short-term excitation pulses with appropriate shapes [32, 33, 34]. Alternatively, one can excite the system by a tightly focused Gaussian beam with a nonzero in-plane momentum, so that pumped polaritons permanently run out of the excitation spot [35, 21, 36, 37]. The third mechanism employs the scattering of the pumped mode into a pair of new macroscopically occupied modes that have different frequencies [38, 39, 40, 41, 42]. In all such situations (see also [43, 44]), some part of a driven system remains free at a certain time, place, or energy level; as a rule, the vortex positions are predetermined. By contrast, our scenario does not rely upon any inhomogeneity of the model and $f$ is constant; at the same time, vortices and domain walls move freely owing to zero-mean $K$. However, this is possible only when $\gamma$ is much smaller than $\Omega$. Relevant parameters of the system of cavity polaritons are discussed in Appendix A.
@@ III Vacuum states
In Appendix B, we find all vacuum states and elementary excitations in a general form. Here, we only illustrate the main features of the relevant solutions at $D\sim\Omega\gg\gamma$.
@@ IV Domain walls in one dimension
In what follows, we solve this boundary-value problem numerically by means of the fourth order collocation algorithm [45] implemented in SciPy [46]. Technically, to do so one has to rewrite (10) in the form of eight real-valued equations of the first order.
@@ IV.1 Static domain walls $(v=0)$
The $q=0$ domain walls [Figs. 3(a) and 3(b)] retain the symmetry of the form $\psi_{+}^{*}=\psi_{-}^{\vphantom{*}}$ for each $x$. Generally, if $\psi_{+}^{*}=\psi_{-}^{\vphantom{*}}$, phases $\phi_{+}$ and $\phi_{-}$ coincide at the center of a domain wall $(x=0)$, being equal to 0 or $\pm\pi$. The solutions with $\phi_{\pm}(0)=0$ never occurred in our calculations; notice that they would have maximized the last term in (2),
$\mathcal{H}_{f}=2f(\sqrt{n_{+}}\cos\phi_{+}+\sqrt{n_{-}}\cos\phi_{-}).$ (14)
In turn, the solutions with $\phi_{\pm}(0)=\pm\pi$ exist in the whole interval $(f_{1},f_{2})$.
@@
Figure 4: Peak values of $n_{\pm}$ (a) and spatial widths $\sigma_{\pm}$ (b) of motionless domain walls depending on $f^{2}$. The dashed vertical line indicates the limiting value of $f^{2}$ at which the $q=\pm 1$ domain walls cease to exist. Parameters as in Fig. 3.
@@ IV.2 Motion of domain walls
Figure 5: Energy (2) (a, b) and peak values of $n_{\pm}$ (c, d) as functions of $v$ for the domain walls with $q=0$ (a, c) and +1+1 (b, d) at $f^{2}/f_{2}^{2}=0.025$. Parameters as in Fig. 3; $H_{\mathrm{vac}}$ is the vacuum energy; $v_{s}=\sqrt{\mu/m}$.
@@
The above predictions have been verified numerically by solving the dynamical equations (21) for a two-dimensional polariton system of size $500\,\xi\times 30\,\xi$. A video presentation of the solution can be found in Supplemental Material I. In the long dimension (axis $x$), the boundary conditions were zero owing to high potential walls at both ends, whereas in the short dimension ($y$) they were periodic. For each $y$, the initial state along the $x$ axis was the $q=1$ domain wall obeying Eqs. (10)–(12) at $D=\Omega$, $\gamma/\Omega=1.5\times 10^{-4}$, and $f_{2}/f_{2}^{2}=0.025$. Finite fluctuations came from the fact that the chosen initial state is not an exact solution for the system in a box. The dynamical equations were solved using the adaptive RK5(4) method by Dormand and Prince [48]. In agreement with our qualitative analysis, it turned out that interaction with fluctuations makes the soliton run in the opposite direction relative to $\nabla\phi_{\pm}$. With time, this soliton becomes “gray” in both spin components, reaches its greatest velocity, and then runs steadily until it hits the end wall. Here, the soliton gets reflected, changes the sign of $q$, and starts moving back towards the other end wall. Thus, the overall dynamics takes the form of periodic spatiotemporal oscillations of $\phi_{\pm}$ and $S_{2}$.
@@
As a final remark, let us notice that the solutions to Eqs. (10) at $v>0$ and $\gamma\to 0$ can also converge to higher-order states with $q\geq 2$. Under normal conditions, such states are unstable in two dimensions, so we do not analyze them in the current work.
@@ V Domain walls in two dimensions
We have found that the $q=\pm 1$ domain walls tend to move in space. Here, we show that they naturally coexist and balance each other in a two-dimensional system. As a characteristic example, Fig. 6 shows a small piece of a domain wall in the vicinity of its crossover between the $q=\pm 1$ states. Two HQVs are seen to connect them continuously, much as a domain wall itself connects phase domains.
The occurrence of “internally ordered” HQV molecules in a driven BEC agrees with the fact that quantity $K$ (5) is small and smooth even at the singular points. To see this, assume that the $n_{+}$ and $n_{-}$ cores lie on the $y=0$ axis at $x=0$ and $x=L>0$, respectively, and the distribution of $\phi_{\pm}$ near the cores corresponds to Figs. 6(c) and 6(d), namely,
$\sin\phi_{+}=-\frac{y}{\sqrt{x^{2}+y^{2}}},\quad\sin\phi_{-}=\frac{y}{\sqrt{(x-L)^{2}+y^{2}}}.$ (16)
In the close vicinity of the $n_{+}$ core, i. e., at $r=\sqrt{x^{2}+y^{2}}\ll L$, condition $K=0$ immediately leads to $\sqrt{n_{+}/n_{-}}=r/L$. Since $n_{-}$ behaves regularly, we finally have $\sqrt{n_{+}}\propto r$ at $r\to 0$, which corresponds to the shape of a conventional vortex. In fact, $K$ is nonzero only on both sides of the $q=\pm 1$ domain walls as shown in Fig. 3(g). Zero mean value of $K$ suggests that the whole pattern represented in Fig. 6 can move freely.
@@ V.1 Synthetic examples
Below, we demonstrate the main kinds of moving collective states with the $q=\pm 1$ junctions. We discuss three solutions to Eqs. (21) obtained on a $1024\times 1024$ grid. In all cases, a high potential wall of radius $R\approx 250\,\xi$ ensures zero boundary conditions. The initial states have two domains placed at $x<0$ and $x>0$ and connected by domain walls $\psi_{\pm}^{(q)}(x)$ whose type $q$ can be different depending on $y$. The driving amplitude $f$ corresponds to Fig. 5 $(f^{2}/f_{2}^{2}=0.025)$. The system dynamics is explicitly shown in the respective Supplemental Materials II, III, and IV that contain video presentations.
@@
Figure 7: Persistently rotating solution with symmetric $q=\pm 1$ segments of a domain wall. (a) and (b) represent phases $\phi_{\pm}$ of two spin components. The arrows indicate the direction of rotation. Parameters: $D=\Omega$, $\gamma/\Omega=7.5\times 10^{-4}$, $f^{2}/f_{2}^{2}=0.025$. The dynamics is shown in Supplemental Material II.
@@
Figure 8: A closed contour of two $q=\pm 1$ domain walls moving in space. (a) and (b) represent phases $\phi_{\pm}$ of two spin components. The arrows indicate the direction of motion. Parameters as in Fig. 7. The dynamics is shown in Supplemental Material III.
@@
Figure 9: Sequence of spatially alternating $q=\pm 1$ states formed after the breakup of the $q=0$ domain wall. (a)–(d) correspond to different time moments. The arrows indicate the directions of motion of small composite solitons. The color map represents $\phi_{+}$. Parameters as in Fig. 7. The dynamics is shown in Supplemental Material IV.
@@
The first example has such initial state that $q=+1$ at $y\geq 0$ and q=1q=-1 at $y<0$. The solution shows that a HQV molecule rapidly appears at $\mathbf{r}=0$, after which the whole picture becomes self-consistent and evolves steadily. As expected, the $q=\pm 1$ parts of the domain wall move in opposite directions and, since they are connected, rotate around $\mathbf{r}=0$. A snapshot of $\phi_{\pm}(\mathbf{r})$ at some time moment is shown in Fig. 7. Clearly, rotation implies that linear velocity $v$ grows with $r$. According to Fig. 5(b), multiple solutions with different $v$ indeed coexist for the same set of boundary conditions. Since the range of $v$ is limited, the uniformly rotating patterns cannot persist against arbitrarily increasing $R$; it is remarkable, however, that they are still possible for $R=250\,\xi$. The stability of the solution is related to the fact that its energy is less than the vacuum one. At the same time, a displacement of the HQV molecule from the center eventually makes the whole system lose its dynamical balance. Analogous patterns often form spontaneously in randomly prepared systems, yet, they usually have smaller sizes and lifetimes.
@@
In the second example, the system evolves from the initial state with $q=+1$ at $|y|\leq a$ and q=1q=-1 at $|y|>a$, where $a$ is a relatively small distance equal to $12.5\,\xi=0.05\,R$. The outer segments have the same $q$ and rotate in opposite directions, so they eventually meet and cancel each other everywhere except the $|y|\lesssim a$ region in which the solution takes the form shown in Fig. 8. It is seen to be composed of two HQV molecules and two $q=\pm 1$ domain walls that encapsulate some piece of one domain inside the other. In spite of opposite $q$, the average phase gradients of the domain walls have the same direction, because the respective pairs of their “left” and “right” domains are also mutually opposite. As a result, we have a composite soliton that moves as a unified object preserving its shape. Direct calculation shows that energy (2) is now greater than the vacuum energy. Thus, being stable in the dynamical sense, this pattern suffers the Landau damping and gradually decreases in size until two HQV molecules annihilate. Noteworthy, it moves in the direction of growing rather than decreasing phases, which corresponds to the bright-soliton section of Fig. 5(b).
@@
The third solution evolves from the most simple initial state: $q=0$ for each $y$. In the considered area of small $f$, such domain walls prove to be dynamically unstable, breaking up in a way similar to the snake (transverse) instability of the $2\pi$ domain walls in atomic BECs that is accompanied by formation of HQV molecules [23, 24]. However, in contrast to atomic BECs, the domain wall as such does not disappear but only reduces its symmetry by turning into a chain of the $q=\pm 1$ states connected by HQV molecules like in Fig. 6. The further evolution is illustrated in Fig. 9. With time, adjacent pairs of molecules annihilate or couple into composite solitons that detach from the wall and decay shortly, producing sound waves. As a result, the total energy of the system decreases.

### v4

@@ Abstract
We consider coherent states of weakly interacting bosons under the conditions of external resonant excitation, with a focus on a two-dimensional polariton fluid driven by a plane electromagnetic wave near the ground state. The coherent driving breaks the U(1) symmetry explicitly, which prevents the occurrence of quantum vortices in a uniform scalar condensate. Surprisingly, a spinor (two-component) system of the same kind admits topological excitations, such as domain walls of relative phase or confined half-vortex molecules, typical of a freely evolving spinor Bose system. Opposite-phase domains arise from the spontaneous breakdown of the spin symmetry $(\mathbb{Z}_{2})$. Domain walls form with time even when the initial state of the system is uniform or completely disordered; they fall into different topological types distinguished by the total phase variation in the transverse direction. One type of domain walls is similar to “magnetic” solitons in Bose-Einstein condensates and exhibits nonzero spin polarization whose sign depends on the direction of motion. Domain walls of the second type, by contrast, behave like monopoles with broken spatiotemporal symmetry and tend to move in certain preferred directions. The interaction of vortices and domain walls results in a long-range ordering of the system.
@@ I Introduction
A coupled pair of spatially overlapping Bose-Einstein condensates (BECs) has complex topological excitations that cannot be reduced to the case of one scalar BEC. Typical examples of such systems include atomic BECs with coexistent hyperfine species [1] and two-band superconductors [2]. In one dimension, they admit the sine-Gordon kinks characterized by the $2\pi$ jump of the relative phase of order parameters [3, 2]; in two dimensions, the kink turns into a domain wall which gives birth to more complicated objects such as bound states of half-quantized vortices (HQVs) [4, 5]. The mechanism of the HQV pairing on a domain wall was found to be formally similar to quark confinement [6, 7]. The underlying mean-field theory attracts much interest owing to very rich physics it contains, in spite of conceptual simplicity, and a broad scope which includes BECs [8], superfluid liquids [9], superconductors [10], and even cosmology [11, 12].
In this work, we investigate a dissipative BEC coupled to an external classical field whose frequency is close to the BEC ground-state level. Such a system is found to admit stable domain walls and HQV molecules which arise spontaneously even when the initial conditions are uniform or, conversely, disordered. A uniform BEC undergoes a spontaneous $\mathbb{Z}_{2}$ symmetry breaking that proceeds nonglobally, by way of the Kibble-Zurek mechanism [13, 14, 15], as soon as the driving field reaches a critical amplitude; a pair of the true (asymptotically stable) vacuum states of the system have opposite phases. On the other hand, in a case of disorder the system exhibits self-organization. Once dynamic equilibrium has established, a number of HQV molecules and domain walls act as elements of complex, inhomogeneous yet internally ordered states whose spatial extent is far beyond any conventional characteristic scale of a BEC.
@@ II Symmetry considerations
The last term in (3) represents the source of excitation. In the case of a one-component system (i. e., when one of $\psi_{\pm}$ is zero or $\psi_{+}$ and $\psi_{-}$ are identical), this driving term invalidates the conventional steady-state vortex solutions such that $\nabla n\neq 0$ and $\nabla\phi\neq 0$ but $(\nabla n)(\nabla\phi)=n\nabla^{2}\phi=0$; moreover, all actual steady states are uniform as long as the driving field has zero wave number [33]. We will show that in the two-component system, by contrast, domain walls and HQV molecules are not only admitted, but arise spontaneously. Domain walls move in space “freely” in the sense that the driving coefficient,
$K=-\sqrt{n_{+}}\sin\phi_{+}-\sqrt{n_{-}}\sin\phi_{-},$ (5)
is zero for each $\mathbf{r}$ for the $q=0$ domain walls and has zero mean for the $q=\pm 1$ ones, provided that $\gamma\to 0$.
The condition $K=0$ is equivalent to $\psi_{+}^{\vphantom{*}}+\psi_{-}^{\vphantom{*}}=\psi_{+}^{*}+\psi_{-}^{*}$ (which represents a sort of weak symmetry) and also equivalent to invariance of the density of energy $(\mathcal{H})$ under the transformation $\psi_{\pm}\mapsto\psi_{\pm}e^{i\chi}$ for small $\chi$. Directly from (2) one can see that $\mathcal{H}\bigl(\psi_{\pm}e^{i\chi}\bigr)-\mathcal{H}\bigl(\psi_{\pm}\bigr)$ tends to $2\chi fK$ when $\chi\to 0$. In fact, it is just a special case of Noether’s theorem that invariance under infinitesimal global phase shift yields the continuity equation with no external sources (e. g., [34]). Therefore, it comes as no surprise that the $q=0$ states with $K=0$ behave similarly to the usual phase domain walls which give birth to HQV molecules via the transverse instability. The symmetry of the $q=\pm 1$ states is reduced further, since $K(\mathbf{r})$ becomes nonzero yet antisymmetric across the wall, which manifests broken parity and involves a row of uncommon features. Nevertheless, when seen as individual objects (on such a scale that their fine structure is indistinguishable), these domain walls also demonstrate an effectively sourceless solitonic motion.
@@
The above considerations help us to compare our results to the previously known mechanisms of vortex formation in resonantly excited polariton systems. In all cases, the key point is to get rid of the phase locking of the driven fluid with respect to the driving optical field. Earlier, this was done by making the system explicitly inhomogeneous in space or time. For instance, vortices can be “injected” directly by means of short-term excitation pulses with appropriate shapes [35, 36, 37]. Alternatively, one can excite the microcavity by a focused laser beam having a nonzero in-plane momentum, so that pumped polaritons run out of the excitation spot and then scatter on artificially introduced impurities [38, 24, 39, 40]. The third mechanism employs the scattering of the pumped mode into a pair of new macroscopically occupied modes that have different frequencies [41, 42, 43, 44, 45]. In all such situations (see also [46, 47]), some part of a driven system gets “unlocked” at a specific place and exhibits topological defects only there. By contrast, our scenario does not rely upon any inhomogeneity of the model and $f$ is constant, so that domain walls and vortices originate spontaneously and move freely. Relevant parameters of the polariton system are discussed in Appendix A.
@@ III Vacuum states
In Appendix B, we find all vacuum states and elementary excitations in a general form. Here, we only illustrate the main features of the relevant solutions at $\gamma\ll D\sim\Omega$.
@@
To simplify the discussion, most of the time we will imply the approximation $\gamma\ll\Omega\sim D$. Let us indicate, however, what changes beyond its scope. In accordance with the exact solutions derived in Appendix B, nonzero $\gamma$ necessarily means nonzero $n_{+}-n_{-}$ and $\alpha_{+}+\alpha_{-}$ (each of these quantities differs in sign for the two vacuum states). Besides, increasing $\gamma$ leads to the increase of the threshold $f_{1}^{2}$ (38) at which the considered vacuum states appear. On the other hand, a significant imbalance of $\Omega$ and $D$ can produce qualitatively different branches of solutions with high $|S_{3}|$ in a range of small $f^{2}$, in which case phase domain walls cease to exist.
@@ IV Domain walls in one dimension
In what follows, we solve this boundary-value problem numerically by means of the fourth-order collocation algorithm [48] implemented in SciPy [49]. Technically, to do so one has to rewrite (10) in the form of eight real-valued equations of the first order.
Figure 4: Peak values of $n_{\pm}$ (a) and spatial widths $\sigma_{\pm}$ (b) of motionless domain walls depending on $f^{2}$. The dashed vertical line indicates the limiting value of $f^{2}$ at which the $q=\pm 1$ domain walls cease to exist. The solution with q=1q=-1 (not shown) is obtained by exchanging ‘++’ and ‘-’ for $q=1$. Parameters as in Fig. 3.
Figure 5: Energy (2) (a, b) and peak values of $n_{\pm}$ (c, d) as functions of $v$ for the domain walls with $q=0$ (a, c) and +1+1 (b, d) at $f^{2}/f_{2}^{2}=0.025$. The solution with q=1q=-1 (not shown) is obtained by reversing time $(v\mapsto-v)$ and exchanging ‘++’ and ‘-’ for $q=1$. Parameters as in Fig. 3; $H_{\mathrm{vac}}$ is the vacuum energy; $v_{s}=\sqrt{\mu/m}$.
@@ IV.1 Static domain walls $(v=0)$
The $q=0$ domain walls [Figs. 3(a) and 3(b)] retain the symmetry of the form $\psi_{+}^{*}=\psi_{-}^{\vphantom{*}}$ for each $x$. Generally, if $\psi_{+}^{*}=\psi_{-}^{\vphantom{*}}$, phases $\phi_{+}$ and $\phi_{-}$ coincide at the center of a domain wall $(x=0)$, being equal to 0 or $\pm\pi$. The solutions with $\phi_{\pm}(0)=0$ never occurred in our calculations; notice that they would have maximized the last term in (2) since $f>0$. In turn, the solutions with $\phi_{\pm}(0)=\pm\pi$ exist in the whole interval $(f_{1},f_{2})$.
@@ IV.2 Motion of domain walls
The above predictions have been verified numerically by solving the dynamical equations (19) for a two-dimensional polariton system of size $500\,\xi\times 30\,\xi$. A video presentation of the solution can be found in Supplemental Material I  [^1] See Supplemental Material I for a video presentation of the soliton motion.. In the long dimension (axis $x$), the boundary conditions were zero owing to high potential walls at both ends, whereas in the short dimension ($y$) they were periodic. For each $y$, the initial state along the $x$ axis was the $q=1$ domain wall obeying Eqs. (10)–(12) at $D=\Omega$, $\gamma/\Omega=1.5\times 10^{-4}$, and $f_{2}/f_{2}^{2}=0.025$. Finite fluctuations came from the fact that the chosen initial state is not an exact solution for the system in a box. The dynamical equations were solved using the adaptive RK5(4) method by Dormand and Prince [52]. In agreement with our qualitative analysis, it turned out that interaction with fluctuations makes the soliton run in the opposite direction relative to $\nabla\phi_{\pm}$. With time, this soliton becomes “gray” in both spin components, reaches its greatest velocity, and then runs steadily until it hits the end wall. Here, the soliton gets reflected, changes the sign of $q$, and starts moving back towards the other end wall. Thus, the overall dynamics takes the form of periodic spatiotemporal oscillations of $\phi_{\pm}$ and $S_{2}$.
@@
Finally, let us notice that the solutions to Eqs. (10) at $v>0$ and $\gamma\to 0$ can also converge to higher-order states with $q\geq 2$. Under normal conditions, such states are unstable in two dimensions, so we do not analyze them in the current work. The greatest positive velocity at which the solutions robustly converge to the $q=1$ states in Figs. 5(b), 5(d) is close to $v_{s}$, where $n_{\pm}$ start growing rapidly. The greatest negative velocity, at which $n_{\pm}$ abruptly drop down, has a smaller absolute value.
@@ V Domain walls in two dimensions
We have found that the $q=\pm 1$ domain walls tend to move in space. Here, we show that they naturally coexist and balance each other in a two-dimensional system. As a characteristic example, Fig. 6 shows a small piece of a domain wall with a crossover between the $q=\pm 1$ states. Two HQVs are seen to connect them continuously, much as a domain wall itself connects phase domains. In the vicinities of both cores, one of the amplitudes $|\psi_{\pm}|=\sqrt{n_{\pm}}$ is proportional to the respective distance. The occurrence of internally ordered HQV molecules in a driven BEC agrees with the fact that quantity $K$ (5) is small and smooth even at the singular points. In fact, $K$ is nonzero only on both sides of the $q=\pm 1$ domain walls in accordance with Fig. 3(g). Zero mean value of $K$ suggests that the whole pattern represented in Fig. 6 can move freely.
@@ V.1 Synthetic examples
Below, we demonstrate the main kinds of moving collective states with the $q=\pm 1$ junctions. We discuss three solutions to Eqs. (19) obtained on a $1024\times 1024$ grid. In all cases, a high potential wall of radius $R\approx 250\,\xi$ ensures zero boundary conditions. The initial states have two domains placed at $x<0$ and $x>0$ and connected by domain walls $\psi_{\pm}^{(q)}(x)$ whose type $q$ can be different depending on $y$. The driving amplitude $f$ corresponds to Fig. 5 $(f^{2}/f_{2}^{2}=0.025)$. The system dynamics is explicitly shown in the respective Supplemental Materials II  [^2] See Supplemental Material II for a video presentation corresponding to Fig. 7., III  [^3] See Supplemental Material III for a video presentation corresponding to Fig. 8., and IV  [^4] See Supplemental Material IV for a video presentation corresponding to Fig. 9. that contain video presentations.
@@
Figure 7: Persistently rotating solution with symmetric $q=\pm 1$ segments of a domain wall. (a) and (b) represent phases $\phi_{\pm}$ of two spin components. The arrows indicate the direction of rotation. Parameters: $D=\Omega$, $\gamma/\Omega=7.5\times 10^{-4}$, $f^{2}/f_{2}^{2}=0.025$. The dynamics is shown in Supplemental Material II [53].
@@
Figure 8: A closed contour of two $q=\pm 1$ domain walls moving in space. (a) and (b) represent phases $\phi_{\pm}$ of two spin components. The arrows indicate the direction of motion. Parameters as in Fig. 7. The dynamics is shown in Supplemental Material III [54].
@@
Figure 9: Sequence of spatially alternating $q=\pm 1$ states formed after the breakup of the $q=0$ domain wall. (a)–(d) correspond to different time moments. The arrows indicate the directions of motion of small composite solitons. The color map represents $\phi_{+}$. Parameters as in Fig. 7. The dynamics is shown in Supplemental Material IV [55].
@@
The first example has such initial state that $q=+1$ at $y\geq 0$ and q=1q=-1 at $y<0$. The solution shows that a HQV molecule rapidly appears near $\mathbf{r}=0$, after which the whole picture becomes self-consistent and evolves steadily. As expected, the $q=\pm 1$ parts of the domain wall move in opposite directions and, since they are connected, rotate around $\mathbf{r}=0$. A snapshot of $\phi_{\pm}(\mathbf{r})$ at some time moment is shown in Fig. 7. Clearly, rotation implies that linear velocity $v$ grows with $r$. According to Fig. 5(b), multiple solutions with different $v$ indeed coexist for the same set of boundary conditions. Since the range of $v$ is limited, the uniformly rotating patterns cannot persist against arbitrarily increasing $R$; it is remarkable, however, that they are still possible for $R=250\,\xi$. The stability of the solution is related to the fact that its energy is less than the vacuum one. At the same time, a displacement of the HQV molecule from the center eventually makes the whole system lose its dynamical balance. Analogous patterns often form spontaneously in randomly prepared systems, yet, they usually have smaller sizes and lifetimes.
@@
In the second example, the system evolves from the initial state with $q=+1$ at $|y|\leq a$ and q=1q=-1 at $|y|>a$, where $a$ is a relatively small distance equal to $12.5\,\xi=0.05\,R$. The outer segments have the same $q$ and rotate in opposite directions, so they eventually meet and cancel each other everywhere except the $|y|\lesssim a$ region in which the solution takes the form shown in Fig. 8. It is seen to be composed of two HQV molecules and two $q=\pm 1$ domain walls that encapsulate some piece of one domain inside the other. In spite of opposite $q$, the average phase gradients of the domain walls have the same direction, because the respective pairs of their “left” and “right” domains are also mutually opposite. As a result, we have a composite soliton that moves as a unified object preserving its shape. Direct calculation shows that energy (2) is now greater than the vacuum energy. Thus, being stable in the dynamical sense, this pattern suffers the Landau damping and gradually decreases in size until two HQV molecules annihilate. Noteworthy, it moves in the direction of growing rather than decreasing phases, which corresponds to the $H>H_{\mathrm{vac}}$ region of Fig. 5(b).
@@
The third solution evolves from the most simple initial state: $q=0$ for each $y$. In the considered area of small $f$, such domain walls prove to be dynamically unstable, breaking up in a way similar to the snake (transverse) instability of the $2\pi$ domain walls in atomic BECs that is accompanied by formation of HQV molecules [26, 27]. However, in contrast to atomic BECs, the domain wall as such does not disappear but only reduces its symmetry by turning into a chain of the $q=\pm 1$ states connected by HQV molecules like in Fig. 6. The further evolution is illustrated in Fig. 9. With time, adjacent pairs of molecules annihilate or couple into composite solitons that detach from the wall and decay shortly, producing sound waves; as a result, the total energy of the system decreases.